State the Domain and Range:
y = 98(-6x - 5)^2
This is the answer I came to:
Domain: all real numbers
Range: all real numbers >=98
I am just wondering if this is correct or if i made a mistake somewhere, thanks for your help!

Answer by ankor@dixie-net.com(12692)

State the Domain and Range:
y = 98(-6x - 5)^2
This is the answer I came to:
Domain: all real numbers
Range: all real numbers >=98
I am just wondering if this is correct or if i made a mistake somewhere, thanks for your help!
I think the range would => 0
:
Check to see if the value inside the brackets can = 0
-6x - 5 = 0
-6x = 5
x = 5/-6
x = -.8333, this value gives y = 0
:
Did this help you?